A probability experiment consists of rolling a 20​-sided die. Find the probability of the event below rolling a number with exactly 3 not necessarily unique prime factors
Answer #1

Let us write the factors of numbers from 1 to 20-

1: 1

2: 1,2

3: 1,3

4: 1,2,4

5: 1,5

6: 1,2,3,6

7: 1,7

8: 1,2,4,8

9: 1,3,9

10: 1,5,10

....

15: 1,3,5,15

15 has 3 unique prime factors- 1 3 and 5

So, we want the probability of the event where a number below 15 is rolled

Required probability = 14/20 = 7/10 = 0.7

Let me know if you need anything else, if not please don't forget to like the answer :)